OpenStudy (anonymous):

I need help with my math practice questions for my test tomorrow, I know how to find the D and R but i hate the one with the questions that contain only x or a square root so if anyone can step by step show me how to find domain and range of y=x^2+1 and y=- square root over x+5

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (triciaal):

y=x^2+1 when x is <0 y is >0 (-x)^2 is positive for any x then y >0 x = set for real numbers -infinity to infinity when x = 0 y = 1 the lowest value of y

OpenStudy (triciaal):

so y (1, infinity)

OpenStudy (triciaal):

y=- square root over x+5 is something missing?

OpenStudy (anonymous):

It says y=-\[\sqrt{x+5}\]

OpenStudy (anonymous):

Plot points

OpenStudy (triciaal):

that's better you had over

OpenStudy (triciaal):

cannot take the rt of a negative (at least for High School) so the minimum value of x is -5

OpenStudy (triciaal):

to plot points pick a value of x and calculate the corresponding value for y eg when x = -5 y =0 when x = 7 y = -rt of 7 + 5 = -rt 12 negative squared is positive and positive squared is positive so for every square root taken there are 2 values the positive and the negative. Here I take it that the value of y is the negative root. need to use one y value for each x value.

OpenStudy (anonymous):

Ok thank you!!!